Solve each equation for x thus expressing x as a log function of y
a) y = 1.3(10)^x
b) y = 8.2(1.3)^x
Can you please help me ? Thank you so much in advance:)

Answer by rothauserc(4718)

from definition of a log
y = b^x, then x = log base b of y
a) y = 1.3(10)^x
y/1.3 = 10^x
x = log (y/1.3)
b) y = 8.2(1.3)^x
y/8.2 = 1.3^x
x = log base 1.3 o (y/8.2)